OpenStudy (anonymous):

Calculate f '(0) for the following continuous function. (If an answer does not exist, enter DNE.) f(x) = x|x|

OpenStudy (zarkon):

use the definition of the derivative

OpenStudy (anonymous):

I'm just not sure how to do that with an absolute value.

OpenStudy (zarkon):

\[f'(0)=\lim_{h\to 0}\frac{f(0+h)-f(0)}{h}\] \[f'(0)=\lim_{h\to 0}\frac{(0+h)|0+h|-0|0|}{h}=\cdots\]

OpenStudy (anonymous):

then it's just 0?

OpenStudy (anonymous):

or does that derivative not exist, because at 0, there's a vertical slope of the tangent line in the function?

OpenStudy (anonymous):

actually, it would be a horizontal slope.

OpenStudy (anonymous):

which would be 0?

OpenStudy (zarkon):

0

OpenStudy (anonymous):

Haha. Ok. Thanks!

OpenStudy (zarkon):

\[f'(0)=\lim_{h\to 0}\frac{(0+h)|0+h|-0|0|}{h}=\cdots\] \[=\lim_{h\to 0}\frac{h|h|}{h}=\lim_{h\to 0}|h|=0\]